OpenStudy (anonymous):

sqrt(4y+5) - sqrt(y-1) = 3 y=??

OpenStudy (anonymous):

not it... I think we are looking for two different values for the y of each side? Perhaps I'm wrong...

OpenStudy (anonymous):

how'd you get 3+(y-1) if you squared the one side it'd be (3+sqroot(y-1))^2

OpenStudy (dumbcow):

koko has the right idea just be careful squaring the right side

OpenStudy (dumbcow):

you should get 3^2 + 2(3sqrt(y-1)) + sqrt(y-1)^2 = 9+6sqrt(y-1) + (y-1) =8 + 6sqrt(y-1) + y

OpenStudy (anonymous):

\[\sqrt{4y+5}=3+\sqrt{y-1}\] square both sides \[4y+5=9+6\sqrt{y-1}+(y-1)\] simplify \[3y-3=6\sqrt{y-1}\] square again \[9y ^{2}-18y+9=36(y-1)\] simplify \[9y ^{2}+18y+45=0\] then factor or use the quadradic

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{4y+5}-\sqrt{y-1}=3\] \[\large \sqrt{4y+5}=3+\sqrt{y-1}\] \[\large 4y+5=(3+\sqrt{y-1})^2\] \[\large 4y+5=9+6\sqrt{y-1}+y-1\] \[\large 4y+5=8+y+6\sqrt{y-1}\] \[\large 4y+5-8-y=6\sqrt{y-1}\] \[\large 3y-3=6\sqrt{y-1}\] \[\large (3y-3)^2=36(y-1)\] \[\large 9y^2-18y+9=36(y-1)\] \[\large 9y^2-18y+9=36y-36\] \[\large 9y^2-18y+9-36y+36=0\] \[\large 9y^2-54y+45=0\] \[\large 9(y^2-6y+5)=0\] \[\large 9(y-1)(y-5)=0\] \[\large y-1=0 \ \ \textrm{or} \ \ y-5=0\] \[\large y=1 \ \ \textrm{or} \ \ y=5\]

OpenStudy (anonymous):

WOW! I'm in over my head!! thnk you:)